# Tight Bounds for Sorting Under Partial Information

### v1

@@ 1 Introduction
Fredman introduced the problem in 1976 [6, TCS’76]. He devised a query strategy that asks $\log e(P)+O(n)$ linear oracle queries. This falls short of the desired $O(\log e(P))$ bound when $\log e(P)$ is sublinear. In 1984, Kahn and Saks [8, Order’84] showed that there always exists a query which reduces the number of linear extensions by a constant fraction; yielding a query strategy that uses $O(\log e(P))$ linear oracle queries. Kahn and Kim [7, STOC’92] propose studying this problem entirely as an algorithmic problem, where in addition to counting the number linear oracle queries used, one also counts the number of algorithmic instructions required to find the next linear oracle query in the query strategy. They note that the algorithm by Kahn and Saks [8] has, in addition to its $O(\log e(P))$ linear oracle queries, exponential running time. This is because in essence their algorithm counts the linear extensions of $P$, which is a #P-complete [1] problem.
@@
$O(n^{2})$ $O(n^{2})$ $O(\log e(P))$ exponential [8, Order’84]
$O(n^{2})$ $O(n^{2})$ $O(\log e(P))$ $O(n^{6})$ [7, STOC’92]
@@
For every constant $c\geq 1$, there is an algorithm that first preprocesses $P$ in $O(n^{1+1/c})$ time, and then asks $O(c\log e(P))$ linear oracle queries in $O(c\log e(P))$ time to recover the linear order on $X$, represented as a leaf-linked tree on $X$.
@@
Let $c$ be a constant greater or equal to $1$. There exists no deterministic algorithm that for all partial orders $P$ can preprocess $P$ in $\frac{1}{48}n^{1+\frac{1}{c}}$ partial oracle queries, such that given the oracle it can recover the linear order on $X$ in at most $\frac{1}{73}c\log e(P)$ exact and partial oracle queries.
@@
There is algorithm that with $\widetilde{O}(n\,w^{2})$ time and queries partitions a poset on $n$ elements into a maximal set $A$ of antichains of size $w+1$, and a set $Y$ of at most $w$ chains.
@@ 2 Preliminaries
A (strict) partial order $P$ on a ground set $X$ is a binary relation on $X$ that is irreflexive, asymetric and transitive. In other words, the digraph on $X$ with an edge $(x_{i},x_{j})$ for every $x_{i}\prec_{P}y_{i}$ is a transitively closed DAG. $(P,X)$ is then called a poset. When clear from context, we omit $X$ and use $P$ to denote the poset $(P,X)$. A linear order (also called total order) is a partial order with no incomparable pairs of elements. The DAG of a linear order contains a Hamilon path. A linear extension $L$ of a partial order $P$ is a linear order $<_{L}$ such that $x_{i}<_{L}x_{j}$ whenever $x_{i}\prec_{P}x_{j}$. A linear extension corresponds to a topological ordering of the DAG $P$. We assume that upper bound algorithms run on a RAM, for a pointer machine algorithm see Appendix A.2.
@@
For output, we recall three previously studied problem variants: Kahn and Saks [8] and Kahn and Kim [7] output a complementing set: a set of directed edges $E\subset X\times X$ where $\forall(x_{i},x_{j})\in E$: $x_{i}<_{L}x_{j}$ and the partial order $P$ (viewed as a DAG) together with $E$ contains a Hamiltonian path. The complementing set has size $O(\log e(P))$ and can therefore be output in $O(\log e(P))$ time. Cardinal [4] have a query algorithm that runs in $O(\log e(P)+n)$ time and may thus output a linked list of $X$ in its sorted order. (HKLS) [12] study the partial order sorting problem for the restricted case where $P$ can be induced from a set of intervals (where two elements are incomparable whenever the intervals overlap). To avoid trivial $\Omega(n)$ lower bounds for reporting a linked list, they propose to output either a linked list or a balanced binary tree on $X$ (in its sorted order), by providing a pointer to either the head of the linked list or the root of the binary tree. Consider the example where $P$ is a linear order and thus $\log e(P)\in O(1)$. This definition of output allows one to preprocess $P$ in $O(n\log n)$ time and output the sorted order of $X$ in $O(1)$ time in the query phase. We preprocess any partial oracle $O_{P}$ to (given $O_{L}$) output either a complementing set, a pointer to a linked list on $X$, or the root of a balanced binary tree on $X$ in its sorted order.
@@
Any deterministic and correct pair of preprocessing and query algorithm on a RAM or pointer machine have a corresponding pair of DTs $(T,\{T_{Q}\})$. The worst-case number of queries performed by the preprocessing algorithm is lower bounded by the height of $T$. The worst-case number of queries performed by the query phase algorithm is lower bounded by $\max_{\ell\in T}\textsc{height}(T_{Q_{\ell}})$.
@@ 3 Algorithm description
During preprocessing, we compute the decomposition of Definition 1. We first remove an (approximately) maximum chain $C_{0}$ from $P$ and store it in a balanced finger search tree $T_{0}$ (in order along $C_{0}$). For some smartly chosen $w$, we compute a maximal set $A$ of antichains of size at least $w+1$ of $P-C_{0}$. This gives a set $Y=P-(C_{0}\cup A)$ which we show we obtain as a set of $w$ chains. We then partition $Y$ into a greedy chain decomposition $C=\{C_{i}^{\prime}\}_{1}^{\ell}$. Each $y\in C$ receives a finger to the maximum $x\in T_{0}$ with $x\prec_{P}y$. each $y\in C$ also receives a finger to the minimal $x\in T_{0}$ with $y\prec_{P}x$.
@@ 4 Analysing the Preprocessing Phase
Consider an algorithm that iteratively considers two elements $y_{1},y_{2}$ that are incomparable in $P$ and removes both of them, until no such pair exists. The remaining set of vertices in $P$ is per definition a chain $C$. We claim that if the longest chain in $P$ has a maximum length of $n-k$ then $C$ has a length of at least $n-2k$. Indeed, let $C^{*}$ be a maximum chain of length $n-k$. Whenever $y_{1},y_{2}$ are incomparable, at least one of them lies outside $C^{*}$. Thus, our procedure deletes at most $k$ pairs of elements, implying $|C|\geq n-2k$.
A straight-forward implementation the above procedure takes $\Theta(n^{2})$ time and partial oracle queries. To speed this up to $O(n\log n)$, we use a sorting algorithm to find incomparable pairs. Indeed, run merge sort on $X$ using the partial oracle $P$. Whenever the merge step encounters two incomparable elements, delete both of them from $X$ and continue with the merge step as usual. When merge sort terminates, the remaining elements are linearly sorted and hence form a chain. ∎
@@
(DKMRV) [5] present an algorithm to partition posets of width $\leq w$ into $w$ chains. In Appendix A, we adapt the core subroutine of of their algorithm (specifically, their ‘peeling’ procedure) to instead iteratively find and delete antichains of size $w+1$. This facilitates a divide-and-conquer approach that arbitrarily splits $P$ into two parts and recursively partitions them into $w$ chains, plus some large antichains. To merge the two parts, we use our new subroutine to extract antichains of size $w+1$. After greedily extracting all large antichains, the remaining elements may be partitioned into $w$ chains by Dilworth’s theorem. ∎
@@ 5 Analysing the Query Phase
Algorithm 2 consists of three components: (1) the merging of $C$ into a chain $C^{*}$, (2) the first for-loop and (3) the second for-loop. We analyse each component separately. For (1), we immediately apply Lemma 4 which runs in $O(|Y|+\log e(P^{\prime}))\subset O(|Y|+\log e(P))$ queries. If $\log e(P)$ is $\log e(P)$ is sublinear, then $|Y|$ is the dominating term. We show that $|Y|\in O(\log e(P))$ and thus this procedure takes $O(\log e(P))$ queries. For (2), we show that merging $C^{*}$ into $C_{0}$ through exponential search uses $O(\log e(P))$ queries. Contrary to previous works that try to efficiently merge chains using the graph entropy, we use a direct counting argument. Finally for (3) we prove Lemma 1: showing that we may insert all elements in the antichains $A$ into $T_{0}$ using binary search in $O(c\log e(P))$ queries, this is the bottleneck of our approach which gives the factor $c$.

### v2

@@ 1 Introduction
Fredman introduced the problem in 1976 [6, TCS’76]. He devised a query strategy that asks $\log e(P)+O(n)$ linear oracle queries. This falls short of the desired $O(\log e(P))$ bound when $\log e(P)$ is sublinear. In 1984, Kahn and Saks [9, Order’84] showed that there always exists a query which reduces the number of linear extensions by a constant fraction; yielding a query strategy that uses $O(\log e(P))$ linear oracle queries. Kahn and Kim [8, STOC’92] propose studying this problem entirely as an algorithmic problem, where in addition to counting the number linear oracle queries used, one also counts the number of algorithmic instructions required to find the next linear oracle query in the query strategy. They note that the algorithm by Kahn and Saks [9] has, in addition to its $O(\log e(P))$ linear oracle queries, exponential running time. This is because in essence their algorithm counts the linear extensions of $P$, which is a #P-complete [1] problem.
@@
$O(n^{2})$ $O(n^{2})$ $O(\log e(P))$ exponential [9, Order’84]
$O(n^{2})$ $O(n^{2})$ $O(\log e(P))$ $O(n^{6})$ [8, STOC’92]
@@
We note that recently and independently, HHIRT [7] also present a subquadratic algorithm for sorting under partial information. Their input is different from their paper, as they assume for input a directed $n$-vertex graph $G$ with $m$ edges which induces a partial order $P$. Their algorithm runs in $O(m+n+\log e(P))$ time and uses $O(\log e(P))$ linear oracle queries and is faster when $m$ is near linear. Their approach cannot separate inspecting the partial order (graph, in their case) and the linear order queries.
@@
For every constant $c\geq 1$, there is a deterministic algorithm that first preprocesses $P$ in $O(n^{1+1/c})$ time, and then asks $O(c\log e(P))$ linear oracle queries in $O(c\log e(P))$ time to recover the linear order on $X$, represented as a leaf-linked tree on $X$.
@@
Let $c$ be a constant greater or equal to $3$. There exists no deterministic algorithm that for all partial orders $P$ can preprocess $P$ in $\frac{1}{48}n^{1+\frac{1}{c}}$ partial oracle queries, such that given the oracle it can recover the linear order on $X$ in at most $\frac{1}{73}c\log e(P)$ exact oracle queries.
@@
There is an algorithm that with $\widetilde{O}(n\,w^{2})$ time and queries partitions a poset on $n$ elements into a maximal set $A$ of antichains of size $w+1$, and a set $Y$ of at most $w$ chains.
@@ 2 Preliminaries
A (strict) partial order $P$ on a ground set $X$ is a binary relation on $X$ that is irreflexive, asymetric and transitive. In other words, the digraph on $X$ with an edge $(x_{i},x_{j})$ for every $x_{i}\prec_{P}y_{i}$ is a transitively closed DAG. $(P,X)$ is then called a poset. When clear from context, we omit $X$ and use $P$ to denote the poset $(P,X)$. A linear order (also called total order) is a partial order with no incomparable pairs of elements. The DAG of a linear order contains a Hamilonian path. A linear extension $L$ of a partial order $P$ is a linear order $<_{L}$ such that $x_{i}<_{L}x_{j}$ whenever $x_{i}\prec_{P}x_{j}$. A linear extension corresponds to a topological ordering of the DAG $P$. We assume that upper bound algorithms run on a RAM, for a pointer machine algorithm see Appendix A.2.
@@
For output, we recall three previously studied problem variants: Kahn and Saks [9] and Kahn and Kim [8] output a complementing set: a set of directed edges $E\subset X\times X$ where $\forall(x_{i},x_{j})\in E$: $x_{i}<_{L}x_{j}$ and the partial order $P$ (viewed as a DAG) together with $E$ contains a Hamiltonian path. The complementing set has size $O(\log e(P))$ and can therefore be outputted in $O(\log e(P))$ time. Cardinal [4] have a query algorithm that runs in $O(\log e(P)+n)$ time and may thus output a linked list of $X$ in its sorted order. (HKLS) [13] study the partial order sorting problem for the restricted case where $P$ can be induced from a set of intervals (where two elements are incomparable whenever the intervals overlap). To avoid trivial $\Omega(n)$ lower bounds for reporting a linked list, they propose to output either a linked list or a balanced binary tree on $X$ (in its sorted order), by providing a pointer to either the head of the linked list or the root of the binary tree. Consider the example where $P$ is a linear order and thus $\log e(P)\in O(1)$. This definition of output allows one to preprocess $P$ in $O(n\log n)$ time and output the sorted order of $X$ in $O(1)$ time in the query phase. We preprocess any partial oracle $O_{P}$ to (given $O_{L}$) output either a complementing set, a pointer to a linked list on $X$, or the root of a balanced binary tree on $X$ in its sorted order.
@@
Any deterministic and correct pair of preprocessing and query algorithm on a RAM or pointer machine has a corresponding pair of DTs $(T,\{T_{Q}\})$. The worst-case number of queries performed by the preprocessing algorithm is lower bounded by the height of $T$. The worst-case number of queries performed by the query phase algorithm is lower bounded by $\max_{\ell\in T}\textsc{height}(T_{Q_{\ell}})$.
@@ 3 Algorithm description
During preprocessing, we compute the decomposition of Definition 1. We first remove an (approximately) maximum chain $C_{0}$ from $P$ and store it in a balanced finger search tree $T_{0}$ (in order along $C_{0}$). For some smartly chosen $w$, we compute a maximal set $A$ of antichains of size at least $w+1$ of $P-C_{0}$. This gives a set $Y=P-(C_{0}\cup A)$ which we obtain as a set of $w$ chains. We then partition $Y$ into a greedy chain decomposition $C=\{C_{i}^{\prime}\}_{1}^{\ell}$. Each $y\in C$ receives a finger to the maximum $x\in T_{0}$ with $x\prec_{P}y$. each $y\in C$ also receives a finger to the minimal $x\in T_{0}$ with $y\prec_{P}x$.
@@ 4 Analysing the Preprocessing Phase
Consider an algorithm that iteratively considers two elements $y_{1},y_{2}$ that are incomparable in $P$ and removes both of them, until no such pair exists. The remaining set of vertices in $P$ is per definition a chain $C$. We claim that if the longest chain in $P$ has a maximum length of $n-k$ then $C$ has a chain of length of at least $n-2k$. Indeed, let $C^{*}$ be a maximum chain of length $n-k$. Whenever $y_{1},y_{2}$ are incomparable, at least one of them lies outside $C^{*}$. Thus, our procedure deletes at most $k$ pairs of elements, implying $|C|\geq n-2k$.
A straight-forward implementation of the above procedure takes $\Theta(n^{2})$ time and partial oracle queries. To speed this up to $O(n\log n)$, we use a sorting algorithm to find incomparable pairs. Indeed, run merge sort on $X$ using the partial oracle $P$. Whenever the merge step encounters two incomparable elements, delete both of them from $X$ and continue with the merge step as usual. When merge sort terminates, the remaining elements are linearly sorted and hence form a chain. ∎
@@
(DKMRV) [5] present an algorithm to partition posets of width $\leq w$ into $w$ chains. In Appendix A, we adapt the core subroutine of their algorithm (specifically, their ‘peeling’ procedure) to instead iteratively find and delete antichains of size $w+1$. This facilitates a divide-and-conquer approach that arbitrarily splits $P$ into two parts and recursively partitions them into $w$ chains, plus some large antichains. To merge the two parts, we use our new subroutine to extract antichains of size $w+1$. After greedily extracting all large antichains, the remaining elements may be partitioned into $w$ chains by Dilworth’s theorem. ∎
@@ 5 Analysing the Query Phase
Algorithm 2 consists of three components: (1) the merging of $C$ into a chain $C^{*}$, (2) the first for-loop and (3) the second for-loop. We analyse each component separately. For (1), we immediately apply Lemma 4 which runs in $O(|Y|+\log e(P^{\prime}))\subset O(|Y|+\log e(P))$ queries. If $\log e(P)$ is sublinear, then $|Y|$ is the dominating term. We show that $|Y|\in O(\log e(P))$ and thus this procedure takes $O(\log e(P))$ queries. For (2), we show that merging $C^{*}$ into $C_{0}$ through exponential search uses $O(\log e(P))$ queries. Contrary to previous works that try to efficiently merge chains using the graph entropy, we use a direct counting argument. Finally for (3) we prove Lemma 1: showing that we may insert all elements in the antichains $A$ into $T_{0}$ using binary search in $O(c\log e(P))$ queries, this is the bottleneck of our approach which gives the factor $c$.